Help! I need assistance trying to find the solution with this problem and need detailed explanations so I can better understand it. I would greatly appreciate it.

Lori, who is single, purchased 5-year class property for $200,000 and 7-year class property for $420,000 on May 20, 2019. Lori expects the taxable income derived from her business (without regard to the amount expensed under § 179) to be about $550,000. Lori has determined she should elect immediate § 179 expensing in the amount of $520,000, but she doesn’t know which asset she should completely expense under § 179. She does not claim any available additional first-year depreciation.

a. Determine Lori’s total deduction if the § 179 expense is first taken with respect to the 5-year class asset.

b. Determine Lori’s total deduction if the § 179 expense is first taken with respect to the 7-year class asset.

c. What is your advice to Lori?

d. Assume that Lori is in the 24% marginal tax bracket and that she elects § 179 on the 7-year asset. Determine the present value of the tax savings from the cost recovery deductions for both assets.

e. Assume the same facts as in part (d), except that Lori decides not to use § 179 on either asset. Determine the present value of the tax savings under this choice.

In addition, determine which option Lori should choose.

Note: please provide answers and explanations on Word or Excel format as written responses are hard to read.
Answer #1

Answer a. Cost recovery deduction if the 179 expense if taken for 5 year class asset

  • Balance deduction for 7 year class asset = 520000 - 5 year asset class value =520000 - 200000=320000
  • balance of 7 year asset class asset after 179 deduction = 420000 - balance =420000-320000=100000
  • 7 year asset class deduction for each year = 100000/7 =14286 (approx)
  • Total deduction available for current year = 520000 +14286 =$ 534286

Answer b. Cost recovery deduction if the 179 expense if taken for 7 year class asset

  • Balane deduction for 5 year class asset under 179 = 520000 - 7 year asset class value =520000 - 420000 =100000
  • 5 year asset class balance after 179 deduction =200000 -100000 =100000
  • 7 year asset class deduction for each year = 100000/5 =20000
  • Total deduction available for current year = 520000 +20000 =$ 540000

Answer c. Advise to lory --------    the best advice to e given to Lori is to move towards higher deduction so that it would result in further deductions as shown below-

option b - section 179 on the 7 year asset. It would result her an additional deduction of (540000-534286=) $ 5714.

Answer d - the present value of the tax savings from the cost recovery deductions for both assets.

please refer to the excel table below -

  • tax deductions/cuttings = 5 year asset + 7 year asset class
  • tax savings = tax deductions * 24 percnt
  • present value/ PV = tax savings * PVF
year 5Year Asset Class 7 Year Asset Class Total cuttings/deductions Tax savings - 24 percnt[GIVEN] PVF PV
1 120000 420000 540000 129600 1.0000 129600
2 32000 0 32000 7680 0.9434 7245.312
3 19200 0 19200 4608 0.8900 4101.12
4 11520 0 11520 2764.8 0.8396 2321.32608
5 11520 0 11520 2764.8 0.7921 2189.99808
6 5760 0 5760 1382.4 0.7473 1033.06752
7 0 0 0 0.7050 0
8 0 0 0 0.6551 0
146490.8237

Answer e - the present value of the tax savings from the cost recovery deductions for both assets, not using 179 expense for either of the assets.

please refer to the excel table below -

  • tax deductions/cuttings = 5 year asset + 7 year asset class
  • tax savings = tax deductions * 24 percnt
  • present value/ PV = tax savings * PVF
year 5Year Asset Class 7 Year Asset Class Total cuttings/deductions Tax savings - 24 percnt[GIVEN] PVF PV
1 40000 60018 100018 24004.32 1.0000 24004.32000
2 64000 102858 166858 40045.92 0.9434 37779.32093
3 38400 73458 111858 26845.92 0.8900 23892.86880
4 23040 52458 75498 18119.52 0.8396 15213.14899
5 23040 37506 60546 14531.04 0.7921 11510.03678
6 11520 37464 48984 11756.16 0.7473 8785.37837
7 37506 37506 9001.44 0.7050 6346.01520
8 18732 18732 4495.68 0.6551 2945.11997
130476.20904
  • higher tax savings would be beneficial. so it recommended to go with option d as it has savings of = 146490.82 - 130476.20 = 16014.61